# On the Curvature of Regge Metrics

### v1

@@ Appendix A Appendix: Nice Manifolds, Blow-Ups, and Integration by Parts
Here we will provide a rapid overview of some key concepts that we have used, many of which are nonstandard. The primary reference for Whitney manifolds (also called “regular manifolds” by Whitney) and Stokes’ Theorem is [19], with [10] providing a gentler introduction. The primary reference for the term “blow-up” as we have used it is [1], although the concept was introduced earlier in at least [14].
@@
A nice $n$-manifold $M$ is a smooth manifold which is locally modeled on relatively open subsets of nondegenerate unions of parallelopipeds. This class of manifolds includes all polytopes and all domains which can be obtained by identifying faces of convex polytopes in $\mathbb{R}^{n}$ by rigid motions.
@@
Specifically, every point $x\in M$ has a coordinate neighborhood $(U,\phi_{U})$ where $\phi_{U}$ is a one-to-one open map $U\to R_{U}\subset\mathbb{R}^{n}$. $R_{U}=P_{1}\cup\dots\cup P_{N}$, where each set $P_{i}$ is a closed nondegenerate $n$-dimensional parallelopiped in $\mathbb{R}^{n}$ and the intersection $P_{i}\cap P_{j}$ is either empty or a shared face of both $P_{i}$ and $P_{j}$. A continuous map $f:A\subseteq R_{U}\to R_{U^{\prime}}$, where $A$ is a relatively open subset of $R_{U}$, is considered a smooth map if it can be extended to a smooth map $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$. The smooth structure is given by a maximal atlas of coordinate patches $(U,\phi_{U})$ such that the transition maps $\phi_{U}\circ\phi_{U^{\prime}}^{-1}$ are all smooth with this definition. As usual, $M$ is also required to be a second-countable Hausdorff space.
@@
Note that if $\psi:A\subseteq R_{U}\to B\subseteq R_{U^{\prime}}$ is a diffeomorphism, then each face $\Delta_{d}\subset\partial R_{U}$ must map to a face $\Delta_{d}^{\prime}$ of $\partial R_{U^{\prime}}$ of the same codimension. For all $0<d\leq n$, let $S_{d}(R_{U})$ be defined as the union of relative interiors of faces of $\partial R_{U}$ which have codimension $d$, and define $S_{0}(R_{U})=\mathring{R}_{U}$. So, if $\phi_{U}(x)\in S_{d}(R_{U})$, then $\phi_{U^{\prime}}(x)\in S_{d}(R_{U^{\prime}})$ for all smooth charts $U^{\prime}\ni x$. The set of such points in $M$ therefore defines a union of disjoint submanifolds (without boundary) $S_{d}(M):=\bigcup_{U}\phi_{U}^{-1}(S_{d}(R_{U}))=\bigcup_{U}\bigcup_{\Delta_{d}\subset S_{d}(R_{U})}\phi_{U}^{-1}(\mathring{\Delta}_{d})$, called the $d$-stratum of $M$. The closure of a connected component of $S_{d}(M)$ is also a nice manifold, and it is called a codimension-$d$ face of $M$, so called because it lies in a codimension-$d$ face in each coordinate chart $U$. Since $\bigcup_{d\geq 1}S_{d}(R_{U})=\partial R_{U}$ for each coordinate chart $U$, we can also define $\partial M:=\bigcup_{d\geq 1}S_{d}(M)$ and $E_{M}:=\bigcup_{d\geq 2}S_{d}(M)$. The set $E_{M}$ is closed and has (n1)(n-1)-dimensional Minkowski content zero in any coordinate chart, and it will be called the exceptional set.
@@
A version of Stokes’ theorem can be produced for nice manifolds, based on the fact that they satisfy all the axioms of Whitney manifolds: if $\omega$ is in $C^{1}\Omega^{n-1}(\mathring{M})$ and bounded on $M\backslash E_{M}$ (meaning each coefficient is bounded in any coordinate chart), $\omega|_{\partial M\backslash E_{M}}$ is summable, and $d\omega|_{\mathring{M}}$ is summable, then [19, p. 108]
@@
What we will call a Blow-Up of a compact nice manifold, called $B_{M}$, is any compact nice manifold of the same dimension as $M$ possessing a continuous onto map of nice manifolds $\Phi:B_{M}\to M$ such that:
@@
    $\Phi|_{\mathring{B}_{M}}:\mathring{B}_{M}\to\mathring{M}$ is a diffeomorphism,
@@
    if $\Delta_{d}$ is a codimension-$d$ face of $M$, $0<d<n$, then $\Phi^{-1}(\mathring{\Delta}_{d})$ is a connected component of $S_{1}(B_{M})$, i.e. it is the relative interior of a codimension-$1$ face of $B_{M}$,
@@
    if $\Delta_{n}$ is a codimension-$n$ face of $M$, then $\Phi^{-1}(\mathring{\Delta}_{n})$ is a codimension-1 face of $B_{M}$ (this distinction is necessary because $\Delta_{n}$ is a single point, so $\mathring{\Delta}_{n}=\Delta_{n}$ is closed),
@@
    $\Phi|_{\Phi^{-1}(\mathring{e})}:\Phi^{-1}(\mathring{e})\to\mathring{e}$ is a diffeomorphism for each codimension-$1$ $e\subset\partial M$,
@@
    $\Phi|_{\Phi^{-1}(\mathring{\Delta}_{d})}:\Phi^{-1}(\mathring{\Delta}_{d})\to\mathring{\Delta}_{d}$ is a smooth map for each codimension-$d$ face $\Delta_{d}\subset\partial M$, $0<d\leq n$,
  6. 6.
[…]
If $M$ is oriented, then we also require that $B_{M}$ is oriented and $\Phi$ is an oriented map.
The easiest example of a nice manifold with a blow-up is the standard unit simplex in $\mathbb{R}^{n}$. Intuitively, the blow-up is the result of intersecting a new half-hyperplane for each $d$-face, $0\leq d<n$, essentially “cutting off” corners, and defining the map $\Phi$ by contracting the newly created faces to the original $d$-faces. The process of cutting off corners is known as omnitruncation, and for a simplex it produces a permutahedron.
Actually writing down an explicit blow-down map seems to be quite difficult; we are aware of [13] where a map from the blown-up simplex (the permutahedron) to the simplex is constructed that has all the needed properties, although only its inverse can be written down explicitly. We also believe that a similar procedure could be used to produce blow-ups of convex polytopes which are also convex polytopes.
@@ Appendix B Appendix: Relationship with Existing Formulas
The following proposition is a rigorous proof that the distributional curvature defined in (22) is equivalent to the equation for the densitized distributional curvature defined by other authors, with a slightly different test space.
@@
Let $f$ be a compatible frame and let $\hat{\phi}$ be a skew-symmetric, compatible $\mathrm{End}(TM)$-valued (n2)(n-2)-form. Then we can define a $(0,4)$-tensor $A(X,Y,Z,W):=(-1)^{n}\left\langle\star\phi(X,Y)W,Z\right\rangle$. Then, using notation from this paper on the left side and notation from [8] on the right side, the following are true:
$\left\langle\hat{R}\wedge\hat{\phi}\right\rangle$ $=\frac{1}{2}\left\langle\mathcal{R},A\right\rangle\omega_{T}$ (28)
$\left\langle\hat{\mathrm{I\!I}}_{e}^{T}\wedge i_{e}^{*}\hat{\phi}^{T}\right\rangle$ $=-2\left\langle\mathrm{I\!I}^{\vec{\nu}^{T}},A_{F\hat{\nu}\hat{\nu}F}\right\rangle\omega_{e}^{T}$ (29)
$\left\langle\hat{\Theta}_{p}\wedge i_{p}^{*}\hat{\phi}\right\rangle$ $=2\Theta_{p}A_{\hat{\mu}\hat{\nu}\hat{\nu}\hat{\mu}}\omega_{p}$ (30)
(To be specific, $\omega_{e}^{T}$ is the induced volume form from the orientation of $T$ on $e$, and $\omega_{p}$ is the induced volume form from the frame $E_{p}$ on $p$.)
@@
To prove part (28) we let $\phi^{i}_{~jkl}:=(-1)^{n}\left\langle\star\hat{\phi}(f_{k},f_{l})f_{j},f_{i}\right\rangle=A(f_{k},f_{l},f_{i},f_{j})$, so $\hat{\phi}=\frac{1}{2}\phi^{i}_{~jkl}f_{i}\otimes f^{j}\otimes\star(f^{k}\wedge f^{l})$, and also write $\hat{R}=\frac{1}{2}\hat{R}(f_{c},f_{d})\otimes(f^{c}\wedge f^{d})$. Note that these are not basis expansions for these expressions, for instance the expression for $\hat{\phi}$ includes a $f_{i}\otimes f^{j}\otimes\star(f^{k}\wedge f^{l})$ term and a $f_{i}\otimes f^{j}\otimes\star(f^{l}\wedge f^{k})$ term for each $i,j,k,l$ (hence the $\frac{1}{2}$ factor). Then we compute:
@@
$=\frac{1}{2}\sum_{i,j,k,l}\left\langle\hat{R}(f_{k},f_{l})f_{i},f_{j}\right\rangle_{T}A(f_{k},f_{l},f_{i},f_{j})\omega_{T}=\frac{1}{2}\left\langle\mathcal{R},A\right\rangle_{T}\omega_{T}.$
To prove part (29), we will use the shorthand $E_{i}^{T}=(E_{e}^{T})_{i}$, where $E_{e}^{T}$ is the orthonormal frame adapted to $e$ using notation from this paper. Without the superscript $T$, $E_{i}$ is implicitly one of the first n1n-1 entries of $E_{e}^{T}$ which do not depend on $T$. Let $\phi^{i}_{~jkl}:=(-1)^{n}\left\langle\star\hat{\phi}^{T}(E_{k}^{T},E_{l}^{T})E_{j}^{T},E_{i}^{T}\right\rangle=A(E_{k}^{T},E_{l}^{T},E_{i}^{T},E_{j}^{T})$ so again $\hat{\phi}=\frac{1}{2}\phi^{i}_{~jkl}E_{i}^{T}\otimes{E^{j}}^{T}\otimes\star({E^{k}}^{T}\wedge{E^{l}}^{T})$. Then we compute:
$\left\langle\hat{\mathrm{I\!I}}_{e}^{T}\wedge i_{e}^{*}\hat{\phi}^{T}\right\rangle=\frac{1}{2}\sum_{k,l,m}\left\langle\hat{\mathrm{I\!I}}_{e}^{T}(E_{m}),\phi^{i}_{~jkl}E_{i}^{T}\otimes{E^{j}}^{T}\right\rangle E^{m}\wedge i_{e}^{*}\star({E^{k}}^{T}\wedge{E^{l}}^{T}).$
@@
$i_{e}^{*}\star({E^{k}}^{T}\wedge{E^{l}}^{T})=\begin{cases}0&\text{if }k,l\neq n,\\ -\star_{e}{E^{k}}^{T}&\text{if }l=n,\\ \star_{e}{E^{l}}^{T}&\text{if }k=n,\end{cases}$
where $\star_{e}$ means the Hodge star operator in the codimension-1 polytope $e$ with the orientation induced by the orientation of $T$. This can be verified using the formula for the Hodge star in an orthonormal coframe along with the fact that the volume form induced on $e$ from the orientation on $T$ is $\omega_{e}^{T}=i_{e}^{*}(E_{n}^{T}\lrcorner\omega_{T})$. Then the previous line simplifies to
$\frac{1}{2}$ $\left\langle\hat{\mathrm{I\!I}}_{e}^{T}(E_{m}),\phi^{i}_{~jkl}E_{i}^{T}\otimes{E^{j}}^{T}\right\rangle E^{m}\wedge i_{e}^{*}\star({E^{k}}^{T}\wedge{E^{l}}^{T})$
@@
$=2\sum_{j,k}\mathrm{I\!I}^{\vec{\nu}^{T}}(E_{k},E_{j})A(E_{k},E_{n}^{T},E_{j},E_{n}^{T})\omega_{e}^{T}=-2\left\langle\mathrm{I\!I}^{\vec{\nu}^{T}},A_{F\hat{\nu}\hat{\nu}F}\right\rangle\omega_{e}^{T}$
@@
To prove part (30), we will use the shorthand $\tau_{i}^{T}=(E_{p,e}^{T})_{i}$, where $E_{p,e}^{T}$ is the positively oriented orthonormal frame adapted to $p$ using notation from this paper. Without the superscript $T$, $\tau_{i}$ is implicitly one of the first n2n-2 entries of $E_{p,e}^{T}$ which do not depend on $e$ or $T$. Let $\phi^{i}_{~jkl}:=(-1)^{n}\left\langle\star\hat{\phi}^{T}(\tau_{k}^{T},\tau_{l}^{T})\tau_{j}^{T},\tau_{i}^{T}\right\rangle=A(\tau_{k}^{T},\tau_{l}^{T},\tau_{i}^{T},\tau_{j}^{T})$, so again $\hat{\phi}=\frac{1}{2}\phi^{i}_{~jkl}\tau_{i}^{T}\otimes{\tau^{j}}^{T}\otimes\star({\tau^{k}}^{T}\wedge{\tau^{l}}^{T})$. Then we compute (abandoning Einstein notation):
@@
$\left\langle\hat{\Theta}_{p}^{T}\wedge i_{p}^{*}\hat{\phi}^{T}\right\rangle$ $=\frac{1}{2}\Theta_{p}\sum_{k,l}\left\langle\tau_{n}^{T}\otimes{\tau^{n-1}}^{T}-\tau_{n-1}^{T}\otimes{\tau^{n}}^{T},\phi^{i}_{~jkl}\tau_{i}^{T}\otimes{\tau^{j}}^{T}\right\rangle i_{p}^{*}\star({\tau^{k}}^{T}\wedge{\tau^{l}}^{T})$
Note that
@@
$i_{p}^{*}\star({\tau^{k}}^{T}\wedge{\tau^{l}}^{T})=\begin{cases}\omega_{p}&\text{if }k=n-1,l=n,\\ -\omega_{p}&\text{if }k=n,l=n-1,\\ 0&\text{otherwise},\end{cases}$
so this simplifies to

### v2

@@ Appendix A Appendix: Polyhedral Manifolds, Blow-Ups, and Integration by Parts
Here we will provide a rapid overview of some key concepts that we have used, some of which are nonstandard. The primary reference for Whitney manifolds (also called “regular manifolds” by Whitney) and Stokes’ Theorem is [25], with [15] providing a gentler introduction. The primary reference for the term “blow-up” as we have used it is [2], although the concept was introduced earlier in at least [19].
@@
A polyhedral $n$-manifold $M$ is a smooth manifold which is locally modeled on relatively open subsets of nondegenerate unions of parallelopipeds. This class of manifolds includes all polytopes and all domains which can be obtained by identifying faces of convex polytopes in $\mathbb{R}^{n}$ by rigid motions.
@@
Specifically, every point $x\in M$ has a coordinate neighborhood $(U,\phi_{U})$ where $\phi_{U}$ is a one-to-one open map $U\to R_{U}\subset\mathbb{R}^{n}$. $R_{U}=P_{1}\cup\dots\cup P_{N_{U}}$, where each set $P_{i}$ is a closed nondegenerate $n$-dimensional parallelopiped in $\mathbb{R}^{n}$ and the intersection $P_{i}\cap P_{j}$ is either empty or a shared face of both $P_{i}$ and $P_{j}$. A continuous map $f:A\subseteq R_{U}\to R_{U^{\prime}}$, where $A$ is a relatively open subset of $R_{U}$, is considered a smooth map if it can be extended to a smooth map $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$. The smooth structure is given by a maximal atlas of coordinate patches $(U,\phi_{U})$ such that the transition maps $\phi_{U}\circ\phi_{U^{\prime}}^{-1}$ are all smooth with this definition. As usual, $M$ is also required to be a second-countable Hausdorff space.
@@
Note that if $\psi:A\subseteq R_{U}\to B\subseteq R_{U^{\prime}}$ is a diffeomorphism, then each face $\Delta_{d}\subset\partial R_{U}$ must map to a face $\Delta_{d}^{\prime}$ of $\partial R_{U^{\prime}}$ of the same codimension. For all $0<d\leq n$, let $S_{d}(R_{U})$ be defined as the union of relative interiors of faces of $\partial R_{U}$ which have codimension $d$, and define $S_{0}(R_{U})=\mathring{R}_{U}$. So, if $\phi_{U}(x)\in S_{d}(R_{U})$, then $\phi_{U^{\prime}}(x)\in S_{d}(R_{U^{\prime}})$ for all smooth charts $U^{\prime}\ni x$. The set of such points in $M$ therefore defines a union of disjoint submanifolds (without boundary) $S_{d}(M):=\bigcup_{U}\phi_{U}^{-1}(S_{d}(R_{U}))=\bigcup_{U}\bigcup_{\Delta_{d}\subset S_{d}(R_{U})}\phi_{U}^{-1}(\mathring{\Delta}_{d})$, called the $d$-stratum of $M$. The closure of a connected component of $S_{d}(M)$ is also a polyhedral manifold, and it is called a codimension-$d$ face of $M$, so called because it lies in a codimension-$d$ face in each coordinate chart $U$. Since $\bigcup_{d\geq 1}S_{d}(R_{U})=\partial R_{U}$ for each coordinate chart $U$, we can also define $\partial M:=\bigcup_{d\geq 1}S_{d}(M)$ and $E_{M}:=\bigcup_{d\geq 2}S_{d}(M)$. The set $E_{M}$ is closed and has (n1)(n-1)-dimensional Minkowski content zero in any coordinate chart, and it will be called the exceptional set.
@@
A version of Stokes’ theorem can be produced for polyhedral manifolds, based on the fact that they satisfy all the axioms of Whitney manifolds: if $\omega$ is in $C^{1}\Omega^{n-1}(\mathring{M})$ and bounded on $M\backslash E_{M}$ (meaning each coefficient is bounded in any coordinate chart), $\omega|_{\partial M\backslash E_{M}}$ is summable, and $d\omega|_{\mathring{M}}$ is summable, then [25, p. 108]
@@
What we will call a Blow-Up of a compact polyhedral manifold, called $B_{M}$, is any compact polyhedral manifold of the same dimension as $M$ possessing a continuous onto map of polyhedral manifolds $\Phi:B_{M}\to M$ such that:
@@
    $\Phi|_{\mathring{B}_{M}}:\mathring{B}_{M}\to\mathring{M}$ is a diffeomorphism.
@@
    For all $d>0$, the preimage of the relative interior of a codimension-$d$ face $\Delta_{d}$ of $M$ is the relative interior of a codimension-$1$ face $\hat{\Delta}_{d}$ of $B_{M}$, and all codimension-$1$ faces of $B_{M}$ are obtained this way.
@@
    $\Phi|_{\Phi^{-1}(\mathring{e})}:\Phi^{-1}(\mathring{e})\to\mathring{e}$ is a diffeomorphism for each codimension-$1$ face $e\subset\partial M$.
@@
    $\Phi|_{\Phi^{-1}(\mathring{\Delta}_{d})}:\Phi^{-1}(\mathring{\Delta}_{d})\to\mathring{\Delta}_{d}$ is a smooth submersion for each codimension-$d$ face $\Delta_{d}\subset\partial M$, $0<d\leq n$.
@@
    Two codimension-1 faces $\hat{\Delta}_{d},\hat{\Delta}_{d^{\prime}}$ of $B_{M}$ have non-empty intersection if and only if $\Delta_{d}\subseteq\Delta_{d^{\prime}}$ or $\Delta_{d^{\prime}}\subseteq\Delta_{d}$.
If $M$ is oriented, then we also require that $B_{M}$ is oriented and $\Phi$ is an orientation-preserving map.
The easiest example of a polyhedral manifold with a blow-up is the standard unit $n$-simplex embedded in $\mathbb{R}^{n}$. Intuitively, the blow-up is the result of intersecting a new half-hyperplane for each $d$-face, $0\leq d<n$, essentially “cutting off” corners, and defining the map $\Phi$ by contracting the newly created faces to the original $d$-faces. The process of cutting off corners is known as omnitruncation, and for a simplex it produces a permutahedron.
Actually writing down an explicit blow-down map between polytopes seems to be quite difficult; we are aware of [18] where a map from the blown-up simplex (the permutahedron) to the simplex is constructed that has all the needed properties, although only its inverse can be written down explicitly. We believe that a similar procedure could be used to produce blow-ups of convex polytopes which are also convex polytopes.
@@ Appendix B Appendix: Relationship with Existing Formulas
The following proposition shows that the distributional curvature defined in (22) is equivalent to the equation for the densitized distributional curvature defined by other authors, with a slightly different test space.
@@
Let $f$ be a compatible frame and let $\hat{\phi}$ be a skew-symmetric, compatible $\mathrm{End}(TM)$-valued (n2)(n-2)-form. Then we can define a $(0,4)$-tensor $A(X,Y,Z,W):=(-1)^{n}\left\langle\star\phi(X,Y)W,Z\right\rangle$. Then, using notation from this paper on the left side and notation from [13] on the right side, the following are true at each point of $\mathring{T}$, $\mathring{e}$, and $\mathring{p}$ respectively:
$\left\langle\hat{R}\wedge\hat{\phi}\right\rangle$ $=\frac{1}{2}\left\langle\mathcal{R},A\right\rangle\omega_{T},$ (28)
$\left\langle\hat{\mathrm{I\!I}}_{e}^{T}\wedge i_{\mathring{e}}^{*}\hat{\phi}|_{T}\right\rangle$ $=-2\left\langle\mathrm{I\!I}^{\vec{\nu}^{T}},A_{F\hat{\nu}\hat{\nu}F}\right\rangle\omega_{e}^{T},$ (29)
$\left\langle\hat{\Theta}_{p}^{T}\wedge i_{\mathring{p}}^{*}\hat{\phi}^{T}\right\rangle$ $=2\Theta_{p}A_{\hat{\mu}\hat{\nu}\hat{\nu}\hat{\mu}}\omega_{p}.$ (30)
In addition, $A_{F\hat{\nu}\hat{\nu}F}$ is well-defined on $\mathring{e}$. To be specific, $\omega_{e}^{T}$ is the induced volume form from the orientation of $T$ on $e$, and $\omega_{p}$ is the induced volume form from the orientation of $e$ on $T$ where $e$ is the side such that $E_{p,e}^{T}$ has an inwards-pointing normal vector.
@@
To prove (28) we let $\phi^{i}_{~jkl}:=(-1)^{n}\left\langle\star\hat{\phi}(f_{k},f_{l})f_{j},f_{i}\right\rangle=A(f_{k},f_{l},f_{i},f_{j})$, so $\hat{\phi}=\frac{1}{2}\phi^{i}_{~jkl}f_{i}\otimes f^{j}\otimes\star(f^{k}\wedge f^{l})$, and also write $\hat{R}=\frac{1}{2}\hat{R}(f_{c},f_{d})\otimes(f^{c}\wedge f^{d})$. Note that these are not basis expansions for these forms, for instance the expression for $\hat{\phi}$ includes a $f_{i}\otimes f^{j}\otimes\star(f^{k}\wedge f^{l})$ term and a $f_{i}\otimes f^{j}\otimes\star(f^{l}\wedge f^{k})$ term for each $i,j,k,l$ (hence the $\frac{1}{2}$ factor). Then we compute:
@@
$=\frac{1}{2}\sum_{i,j,k,l}\left\langle\hat{R}(f_{k},f_{l})f_{i},f_{j}\right\rangle_{T}A(f_{k},f_{l},f_{i},f_{j})\omega_{T}=\frac{1}{2}\left\langle\mathcal{R},A\right\rangle\omega_{T}.$
To prove (29), we will use the shorthand $E_{i}^{T}=(E_{e}^{T})_{i}$, where $E_{e}^{T}$ is the orthonormal frame adapted to $e$ using notation from this paper. Without the superscript $T$, $E_{i}$ is implicitly one of the first n1n-1 entries of $E_{e}^{T}$ which do not depend on $T$. Let $\phi^{i}_{~jkl}:=(-1)^{n}\left\langle\star_{T}\hat{\phi}|_{T}(E_{k}^{T},E_{l}^{T})E_{j}^{T},E_{i}^{T}\right\rangle_{T}=A|_{T}(E_{k}^{T},E_{l}^{T},E_{i}^{T},E_{j}^{T})$ so again $\hat{\phi}=\frac{1}{2}\phi^{i}_{~jkl}E_{i}^{T}\otimes{E^{j}}^{T}\otimes\star_{T}({E^{k}}^{T}\wedge{E^{l}}^{T})$. Then we compute:
$\left\langle\hat{\mathrm{I\!I}}_{e}^{T}\wedge i_{\mathring{e}}^{*}\hat{\phi}|_{T}\right\rangle=\frac{1}{2}\left\langle\hat{\mathrm{I\!I}}_{e}^{T}(E_{m}),\phi^{i}_{~jkl}E_{i}^{T}\otimes{E^{j}}^{T}\right\rangle E^{m}\wedge i_{\mathring{e}}^{*}\star_{T}({E^{k}}^{T}\wedge{E^{l}}^{T}).$
@@
$i_{\mathring{e}}^{*}\star_{T}({E^{k}}^{T}\wedge{E^{l}}^{T})=\begin{cases}0&\text{if }k,l\neq n,\\ -\star_{e}{E^{k}}^{T}&\text{if }l=n,k\neq n\\ \star_{e}{E^{l}}^{T}&\text{if }k=n,l\neq n\end{cases}$
where $\star_{e}$ means the Hodge star operator in the codimension-1 polytope $e$ with the orientation induced by the orientation of $T$. This can be verified using the formula for the Hodge star in an orthonormal coframe along with the fact that the volume form induced on $e$ from the orientation on $T$ is $\omega_{e}^{T}=i_{\mathring{e}}^{*}(E_{n}^{T}\lrcorner\omega_{T})$. Then the previous line simplifies to
$\frac{1}{2}$ $\left\langle\hat{\mathrm{I\!I}}_{e}^{T}(E_{m}),\phi^{i}_{~jkl}E_{i}^{T}\otimes{E^{j}}^{T}\right\rangle E^{m}\wedge i_{\mathring{e}}^{*}\star_{T}({E^{k}}^{T}\wedge{E^{l}}^{T})$
@@
$=2\sum_{j,k}\mathrm{I\!I}^{\vec{\nu}^{T}}(E_{k},E_{j})A|_{T}(E_{k},E_{n}^{T},E_{j},E_{n}^{T})\omega_{e}^{T}=-2\left\langle\mathrm{I\!I}^{\vec{\nu}^{T}},A_{F\hat{\nu}\hat{\nu}F}\right\rangle\omega_{e}^{T}.$
@@
We still need to show that $A_{F\hat{\nu}\hat{\nu}F}$ is well-defined. Assume that $T$ is the side of $e=T\cap T^{\prime}$ such that $E_{e}^{T}$ is positively oriented. Then we have
$A_{F\hat{\nu}\hat{\nu}F}(E_{i},E_{j})|_{T}$ $=(-1)^{n}\left\langle\star_{T}\hat{\phi}|_{T}(E_{i},E_{n}^{T})E_{j},E_{n}^{T}\right\rangle_{T}$
$=(-1)^{i-1}\left\langle\hat{\phi}|_{T}(E_{1},\dots,\hat{E}_{i},\dots E_{n-1})E_{j},E_{n}^{T}\right\rangle_{T}$
$=(-1)^{i-1}{[i_{\mathring{e}}^{*}\hat{\phi}|_{T}(E_{1},\dots,\hat{E}_{i},\dots,E_{n-1})]_{E_{e}^{T}}}^{n}_{j}$
$=(-1)^{i-1}\mathrm{Ad}\left(\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\right)\left([i_{\mathring{e}}^{*}\hat{\phi}|_{T^{\prime}}(E_{1},\dots,\hat{E}_{i},\dots,E_{n-1})]_{E_{e}^{T^{\prime}}}\right)^{n}_{j}$
$=(-1)^{i}{[i_{\mathring{e}}^{*}\hat{\phi}|_{T^{\prime}}(E_{1},\dots,\hat{E}_{i},\dots,E_{n-1})]_{E_{e}^{T^{\prime}}}}^{n}_{j}$
$=(-1)^{i}\left\langle\hat{\phi}|_{T^{\prime}}(E_{1},\dots,\hat{E}_{i},\dots,E_{n-1})E_{j},E_{n}^{T^{\prime}}\right\rangle_{T^{\prime}}.$
The notation $\hat{E}_{i}$ above means the $i$th vector is skipped. Here we need to apply the fact that the frame $E_{e}^{T^{\prime}}$ is negatively oriented on $M$, so that $\star_{T^{\prime}}\hat{\phi}|_{T^{\prime}}(E_{i},E_{n}^{T^{\prime}})=(-1)^{n+i}\hat{\phi}|_{T^{\prime}}(E_{1},\dots,\hat{E}_{i},\dots,E_{n-1})$. This shows that $A_{F\hat{\nu}\hat{\nu}F}|_{T^{\prime}}=A_{F\hat{\nu}\hat{\nu}F}|_{T}$.
@@
To prove (30), we will use the shorthand $\tau_{i}^{T}=(E_{p,e}^{T})_{i}$, where $E_{p,e}^{T}$ is the positively oriented orthonormal frame adapted to $p$ using notation from this paper, and $e\supset p$ is the facet for which $\tau_{n}=-\vec{n}^{T}$. Without the superscript $T$, $\tau_{i}$ is implicitly one of the first n2n-2 entries of $E_{p,e}^{T}$ which do not depend on $e$ or $T$. Let $\phi^{i}_{~jkl}:=(-1)^{n}\left\langle\star\hat{\phi}|_{T,e}(\tau_{k}^{T},\tau_{l}^{T})\tau_{j}^{T},\tau_{i}^{T}\right\rangle_{T}=A|_{T,e}(\tau_{k}^{T},\tau_{l}^{T},\tau_{i}^{T},\tau_{j}^{T})$, so again $\hat{\phi}|_{T,e}=\frac{1}{2}\phi^{i}_{~jkl}\tau_{i}^{T}\otimes{\tau^{j}}^{T}\otimes\star({\tau^{k}}^{T}\wedge{\tau^{l}}^{T})$.
Here, because $\hat{\phi}|_{T}$ and $A|_{T}$ are both discontinuous on $\mathring{p}$, care needs to be taken about which of the “representatives” we use to evaluate $\phi^{i}_{~jkl}$. We used the notation $\hat{\phi}|_{T,e}$ to refer to the value of $\hat{\phi}|_{T}$ that is continuously extended from $\mathring{e}$, i.e. $\hat{\phi}|_{T,e}(x)=\lim_{m\to\infty}\hat{\phi}|_{T}(x_{m})$ where $x_{m}$ is any sequence of points in $\mathring{e}$ that converge to $x\in p$, and likewise for $A$. This choice means that $i_{\mathring{p}}^{*}\hat{\phi}|_{T,e}={\chi_{0,p}^{T}}^{*}\hat{\phi}^{T}$, although as discussed at the end of subsection 2.5, the product we are evaluating ultimately does not depend on this choice.
Then we compute (abandoning Einstein notation):
@@
$\left\langle\hat{\Theta}_{p}^{T}\wedge{\chi_{0,p}^{T}}^{*}\hat{\phi}^{T}\right\rangle$ $=\frac{1}{2}\Theta_{p}\sum_{i,j,k,l}\left\langle\tau_{n}^{T}\otimes{\tau^{n-1}}^{T}-\tau_{n-1}^{T}\otimes{\tau^{n}}^{T},\phi^{i}_{~jkl}\tau_{i}^{T}\otimes{\tau^{j}}^{T}\right\rangle i_{\mathring{p}}^{*}\star({\tau^{k}}^{T}\wedge{\tau^{l}}^{T})$
Note that, because $\tau_{n}^{T}=-\vec{n}^{T}$,
@@
$i_{\mathring{p}}^{*}\star({\tau^{k}}^{T}\wedge{\tau^{l}}^{T})=\begin{cases}-\omega_{p}&\text{if }k=n-1,l=n,\\ \omega_{p}&\text{if }k=n,l=n-1,\\ 0&\text{otherwise},\end{cases}$
where $\omega_{p}=-\tau^{1}\wedge\dots\wedge\tau^{n-2}$ is the volume form induced on $p$ by the orientation of $e$. So this simplifies to